# Local Attention Mechanism: Boosting the Transformer Architecture for Long-Sequence Time Series Forecasting

### v2

@@ Abstract
Keywords Disease detection  $\cdot$ Deep Learning  $\cdot$ Small Data  $\cdot$ Computer Vision
@@ 1 Introduction
We are currently living in the information age, with vast amounts of data being continuously generated and requiring processing to maximise their utility. From this type of data, time series emerge as the natural choice for representation in several applications [1]. Time series, which consist of sequentially related data, are commonly encountered in sectors such as finance [2], industry [3], telecommunications [4], electricity [5, 6], and climatology [7, 8]. Time series forecasting involves predicting future time instances based on past information. An specific scenario of this problem is Long Sequence Time Series Forecasting (LSTF) [9], which implies a longer prediction horizon and represents the most interesting and complex problem in time series forecasting. One of the main concerns of the LSTF problem is capturing long-term trends in the data. To adequately predict these long-term patterns, a substantial number of instances is required to model complex behaviour effectively.
@@
Traditionally, the forecasting problem has been addressed using statistical methods like ARIMA [10]. However, these models are prone to significant long-term error accumulation [11], prompting the exploration of new alternatives within the field of Deep Learning. Convolutional models for time series [12, 13] leverage the concept correlation in consecutive time steps. Various models have been developed based on this architecture, typically incorporating recurrent layers [14] such as Long-Short Term Memory (LSTM) [15, 16, 17]. This type of network faces clear limitations due to the complexity of recurrent operations [18] and the lack of capacity to capture long-term time dependencies [19]. Transformers, originally proposed for natural language processing [20], have been incorporated into the field of time series and forecasting to address these needs. Nonetheless, current transformer models present two main limitations. First of all, the computational evaluation of the attention mechanism requires significant time and memory due to its quadratic complexity on the length of the input. Secondly, the utilisation of attention mechanisms results in a significant loss of positional information, which poses a critical challenge for time series forecasting [21]. As a first step to address the computational cost issue, an efficient attention mechanism based on probabilistic sampling of attentions has been proposed in [22]. However, to date, methodologies employing transformers for time-series forecasting have not sufficiently mitigated the loss of positional information [21]. Consequently, emerges a necessity for an innovative attention mechanism that leverages the local and long-term temporal dependency properties of time series, does not incur information loss and minimises computational and memory overheads.
@@
Previously existing transformer models within the field of time series do not exploit this conception of locality and present experimental studies with data sets of limited size, not only in terms of features but particularly in terms of instances. If the goal is capturing long-term dependencies, large datasets are needed to adequately train the model weights. The primary defining feature of a time series is the dependence of past time steps on the current time instant [23]. This dependence is more likely to be higher in closer time steps, specially in the case where the time series comes from discretizising a continuous stochastic process.
@@
To validate the proposed approach, the second objective outlined above will be pursued to generate a robust dataset for performance analysis. For fairness and consistency, we will first utilize the datasets employed in the Informer papers and other state-of-the-art methods as an initial benchmark, before incorporating our own proposed dataset into the study.
The rest of this paper is structured as follows. In Section 2, we discuss the background on transformers and the LSTF problem. In Section 3, we detail our proposed mechanism, introducing all relevant details. In Section 4, we compare our approach with the state-of-the-art methods using currently employed LSTF datasets. In Section 5, we extend the datasets used to better represent the complexities of the LSTF problem and establish a more comprehensive benchmark. Finally, we present our conclusions and outline future work in Section 6.
@@ 2 Transformers for Long Sequence Time-series Forecasting
A time series is a sequence of random vectors $X_{t}\in\mathbb{R}^{d}$ recorded at regular time intervals, typically denoted as $\{X_{t}\}_{t=1}^{T}$. We use $X_{t}$ to refer to the random vector and $x_{t}$ to represent the observation at time $t$ [24]. Time series data is characterised by its temporal dependency, where each observation relies on previous values, potentially exhibiting long-term patterns [25]. Such data often features distinct elements: trend (long-term data direction), seasonality (predictable periodic fluctuations) and cyclic patterns (less regular variations influenced by external factors). In this work we focus on time series forecasting, which corresponds to the problem of estimating $\mathbb{E}\left[\left.X_{t+1}\right|X_{1}=x_{1},\ldots,X_{t}=x_{t}\right]$. Specifically, we address long sequence time-series forecasting (LSTF)[9], where the goal is predicting long sequences of future data points in datasets with extensive historical records, requiring of the application of advanced techniques to understand complex patterns and trends.
Neural networks, particularly one-dimensional convolutional neural networks (1D CNNs) and Long Short-Term Memory (LSTM) networks, offer black-box solutions for time-series forecasting [26, 27]. 1D CNNs use convolutional layers to identify local patterns and hierarchies within time series data, but are unable to capture longer time dependencies [19]. Indeed, short-term prediction are greatly affected by locality, that is, the last few elements of the sequence greatly determine the trend of the model and, thus, the short-term prediction, unless seasonality and cyclic patterns play a role at that specific timestamp. LSTMs are able to take into account seasonality and cyclic patterns for medium-term dependencies thanks to the use memory cells, overcoming the vanishing gradient problem via gating mechanisms. However, the recursive nature of LSTMs and the memory cells limitations makes them less efficient and effective for long-term predictions [22].
Transformers have surpassed RNNs and LSTMs in several applications, such as NLP, thanks to their ability to model long-range dependencies, learn from large datasets, and parallelise computations efficiently. Motivated by these properties, transformers have been deemed as excellent candidates to the LSTF problem [28, 22]. Indeed, unlike traditional methods that rely on fixed time windows, lagged observations and/or recursive neural networks, transformers can process long sequences efficiently, making them suitable for both short-term and long-term forecasting tasks [29].
Transformers are complex encoder-decoder architectures that encode the sequence of samples and use the extracted information to predict the several future elements of the sequence via a decoder. The key element to understand these architectures is the attention mechanism, see [20]. In a nutshell, the main conclusion of [20] is that performance in NLP applications can be significantly improved by eliminating the recurrence structure of deep learning models and instead focusing exclusively on exploiting the attention mechanism to extract relevant information from sequences of data. In this work we refer to this encoding-decoding model as the vanilla transformer.
When it comes to the application of transformers to time-series problems, progress has been significantly slower than in NLP. This is due to the fact that the attention mechanism presents two main limitations: loss of positional information, and computational and memory cost [28]. The former problem is solved in the application of the vanilla transformer to NLP problems via a positional encoding of the input, and residual connections, which we will be explained in Section 2.3 and Section 2.4 respectively. However, the validity of this approach in the case of time series remains controversial, see, for instance [21], suggesting that further research is needed. We will come back to this problem when introducing our attention mechanism, LAM, in Section 3, which greatly reduces the impact of loss of positional information. On the other hand, improving the computational and memory cost of the attention mechanism has actually been addressed in the time-series literature. The evaluation of the original attention mechanism has time and memory complexity $\Theta(n^{2})$, where $n$ is the width of this layer. As a consequence of the memory bottleneck, not every data scientist has access to the computing resources needed to train and evaluate a transformer model with the full attention mechanism on large sequences of data, preventing transformers from becoming a machine learning state of the art in time series forecasting in practice. The most remarkable solution to accelerate the attention mechanism for time series forecasting is that of [22], where the authors propose a probabilistic attention mechanism that runs in time and memory $\Theta(n\log n)$. Other proposed solutions to this problem involve applying other sorts of mechanisms to the sequence of data instead of the attention mechanism, see for instance [30].
@@
In the rest of this section, we describe the elements of the vanilla transformer, which will serve as the base model for our experimentation, and the probabilistic attention mechanism of [22]. This section is organised as follows. In Section 2.1 we briefly recapitulate the canonical attention mechanism introduced in [20]. In Section 2.2 we summarise the concept of probabilistic attention and point out some of its weaknesses. In Section 2.3 we describe the positional encoding used in our experiments. Finally, in Section 2.4 we summarise the encoder-architecture of the vanilla transformer.
@@ 2.1 Attention Mechanism
so the $i$-th output is a weighted average of the values. The scale factor $1/\sqrt{d_{q}}$ does not change the output of $\operatorname{softmax}$; its purpose is to help scaling the gradients of $\operatorname{softmax}$ during training in order to avoid gradients too close to 00, which may decrease the rate at which a transformer is able to learn from data.
The most common application of attention in transformers corresponds to the case $Q=K$. In this setting, the full attention mechanism resembles a smooth non-parametric regression [31, 32] of $V$ in terms of $Q=K$,  [33], effectively providing a smoothing of $V$. We refer to self-attention as the case when $Q=K=V$. Intuitively, self-attention weighs in the similarity of different elements in a sequence and smooths the input data. Empirically this process has been shown to be highly effective in capturing both local and global patterns in data [34]. However, this intuition has not been corroborated from a theoretical point of view, where mathematical results have only managed to unveil the Lipschitz constant of the attention mechanism [35].
@@
where $\text{where }head_{i}$ $=\operatorname{Attn}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),$
@@
A critique to the full attention mechanism is the potential loss of positional information. First of all, note that the output of $\operatorname{FullAttn(Q,K,V)}$ is a matrix whose $i$-th row is $\sum_{j=1}a_{i,j}V_{j}$. This average $\sum_{j=1}a_{i,j}V_{j}$ results in a partial loss of position as it may be the case that $a_{i,j}$ is large for a $V_{j}$ that is far away from $V_{i}$ in the input sequence. That is, we lose the sense of temporal locality of the values as we average values with entries far away in the input sequence. We formalise this loss of local information is in Proposition 1. First, let us introduce some notation. For a permutation $\pi\colon\{0,1,\ldots,n-1\}\to\{0,1,\ldots,n-1\}$, the permutation matrix $P_{\pi}$ is the matrix with dimensions $n\times n$ defined as $P_{\pi}[i,j]=1$ for $j=\pi(i)$ and $P_{\pi}[i,j]=0$ otherwise.
@@
Proof.
We first prove the proposition for $h=1$, that is, we show that
As we will see in Section 2.4, usually only the last layer of the vanilla transformer is a fully-linear layer; the rest of layers are attention layers with projections acting on the dimension of the characteristics. Therefore, positional information is only exploited in these last fully-linear layers. However, during the repeated application of attention layers positional information is reduced due to the averages carried out in the full attention mechanism. This issue is addressed by the implementation of the positional encoding, which aims to minimise this reduction of positional information when applying attention layers. However, it seems that, in the case of time series, positional encoding is not enough. Indeed, in experiments, on average a vanilla transformer trained for a particular time series produces similar predictions on an original input and a permuted version of this input  [21].
@@ 2.2 Probabilistic Attention
Among state-of-the-art models, Informer is distinguished as one of the most frequently cited and top-performing [22]. The authors propose a probabilistic attention mechanism that runs in time and memory $\Theta(n\log n)$. The key idea here is selecting $u$ queries, where $u=5\lceil\log n\rceil$, and setting the rest of the queries to zeroes. To select these queries, the authors introduce a sparsity measurement for the queries, which approximates the Kullback-Leibler divergence of the attention scores of a query and the uniform distribution. Finding the queries whose attention scores have largest sparsity measurement would require to compute the whole attention matrix to begin with. Instead, the authors propose a heuristic method: for each query $Q_{i}$, they sample $\Theta(\log n)$ scalar products of the form $\langle Q_{i},K_{j}\rangle$ and approximate the sparsity measurement of the query $Q_{i}$ using these samples. Let us denote by $\operatorname{ProbSelect}(Q,K)$ the method that selects the top $u$ queries of $Q$ with this sampling heuristic and masks the rest of the queries to zeroes. Then, the probabilistic attention mechanism is
@@
Even though the ProbAttn method does significantly reduce the time and memory complexity of the attention mechanism and seems to be able to capture the most significant attention scores in practice according to the experiments carried out in [22], it does lead to a significant loss of information. Indeed, the output of $\operatorname{ProbAttn(Q,K,V)}$ has $n-u$ entries that are equal to the average of the values. We believe this to be a major drawback of this attention mechanism, and this is corroborated in our experimentation. The authors of [22] try to fix this issue by applying residual connections in their model, as well as a complex architecture that exploits convolutions in conjunction with ProbAttn. However, our experiments show that this is not enough, as the vanilla transformer is able to significantly outperform Informer.
2.3 Positional encoding
2.4 The Encoder-Decoder architecture
The encoder-decoder architecture in transformers forms the backbone of many advanced models in natural language processing and beyond. In this architecture, the encoder processes the input sequence, transforming it into a rich, contextual representation. Here we briefly describe the architecture proposed in [20], which we use in our model in conjunction with LAM. Each encoder layer consists of a multi-head self-attention mechanism, which allows the model to weigh the importance of different parts of the input relative to each other, and two linear projections acting on the dimension of the variables (i.e. the last dimension of the input) with activation functions. Formally, these linear projections act as the matrix $W^{O}$ in the definition of $\operatorname{MultiHead}$, they are represented as $d_{model}\times d_{model}$ matrices, but they have LeakyReLu activation functions on the outputs. The purpose of projection layers is to learn a domain representation that boosts the performance of the attention mechanism.

### v3

@@ Abstract
Keywords time series, transformer, attention mechanism, long sequence time series forecasting, time series forecasting
@@ 1 Introduction
We are currently living in the information age, with vast amounts of data being continuously generated and requiring processing to maximise their utility. From this type of data, time series emerge as the natural choice for representation in several applications (lin2003symbolic, ). Time series, which consist of sequentially related data, are commonly encountered in sectors such as finance (zhang2024deep, ), industry (yan2024comprehensive, ), telecommunications (di2024hybrid, ), electricity (gulay2024forecasting, ), and climatology (karevan2020transductive, ).
Time series forecasting involves predicting future time instances based on past information. An specific scenario of this problem is Long Sequence Time Series Forecasting (LSTF) (chen2023long, ), which implies a longer prediction horizon and represents the most interesting and complex problem in time series forecasting. One of the main concerns of the LSTF problem is capturing long-term trends in the data. To adequately predict these long-term patterns, a substantial number of instances is required to model complex behaviour effectively.
@@
Traditionally, the forecasting problem has been addressed using statistical methods like ARIMA (arima, ). However, these models are prone to significant long-term error accumulation (arima-limitations, ), prompting the exploration of new alternatives within the field of Deep Learning. Convolutional models for time series (8489399, ; electronics8080876, ) leverage the concept correlation in consecutive time steps. Various models have been developed based on this architecture, typically incorporating recurrent layers (HEWAMALAGE2021388, ) such as Long-Short Term Memory (LSTM) (10.1115/1.4056122, ; DUDUKCU2023109945, ; AGUILERAMARTOS2023223, ). This type of network faces clear limitations due to the complexity of recurrent operations (lipton2015critical, ) and the lack of capacity to capture long-term time dependencies (al2023lstm, ).
Transformers, originally proposed for natural language processing (vaswani2017attention, ), have been incorporated into the field of time series and forecasting to address these needs. Nonetheless, current transformer models present two main limitations. First of all, the computational evaluation of the attention mechanism requires significant time and memory due to its quadratic complexity on the length of the input. Secondly, the utilisation of attention mechanisms results in a significant loss of positional information, which poses a critical challenge for time series forecasting (zeng2023transformers, ). As a first step to address the computational cost issue, an efficient attention mechanism based on probabilistic sampling of attentions has been proposed in (zhou2023informer, ). However, to date, methodologies employing transformers for time-series forecasting have not sufficiently mitigated the loss of positional information (zeng2023transformers, ). Consequently, emerges a necessity for an innovative attention mechanism that leverages the local and long-term temporal dependency properties of time series, does not incur information loss and minimises computational and memory overheads.
@@
Previously existing transformer models within the field of time series do not exploit this conception of locality and present experimental studies with data sets of limited size, not only in terms of features but particularly in terms of instances. If the goal is capturing long-term dependencies, large datasets are needed to adequately train the model weights. The primary defining feature of a time series is the dependence of past time steps on the current time instant (bloomfield2004fourier, ). This dependence is more likely to be higher in closer time steps, specially in the case where the time series comes from discretizising a continuous stochastic process.
@@
To validate the proposed approach, we have divided the experimental framework in three scenarios. First, LAM will be compared within the state-of-the-art in transformer models and most used datasets. Secondly, Informer’s attention mechanism (ProbAttention) and LAM will be compared under the same architecture to provide a fair comparison. Finally, we propose a new dataset collection for the LSTF problem, requiring a wider prediction horizon and larger datasets. Thanks to this comprehensive comparison, we cover all possible cases and provide robust results.
The rest of this paper is structured as follows. In Section 2, we discuss the background on transformers and the LSTF problem. In Section 3, we detail LAM  introducing all relevant details. In Section 4, we compare our approach with the state-of-the-art methods using currently employed LSTF datasets. In Section 5, we compare ProbAttention and LAM within the vanilla transformer architecture. In Section 6, we extend the datasets used to better represent the complexities of the LSTF problem and establish a more comprehensive benchmark. Finally, we present our conclusions and outline future work in Section 7.
@@ 2 Transformers for Long Sequence Time-series Forecasting
A time series is a sequence of random vectors $X_{t}\in\mathbb{R}^{d}$ recorded at regular time intervals, typically denoted as $\{X_{t}\}_{t=1}^{T}$. We use $X_{t}$ to refer to the random vector and $x_{t}$ to represent the observation at time $t$ (lai2018modeling, ). Time series data is characterised by its temporal dependency, where each observation relies on previous values, potentially exhibiting long-term patterns (hyndman2018forecasting, ). Such data often features distinct elements: trend (long-term data direction), seasonality (predictable periodic fluctuations) and cyclic patterns (less regular variations influenced by external factors). In this work we focus on time series forecasting, which corresponds to the problem of estimating $\mathbb{E}[X_{t+1}|X_{1}=x_{1},\ldots,X_{t}=x_{t}]$. Specifically, we address LSTF problem (chen2023long, ), where the goal is predicting long sequences of future data points in datasets with extensive historical records, requiring of the application of advanced techniques to understand complex patterns and trends.
@@
In the rest of this section, we describe the elements of the vanilla transformer, which will serve as the base model for our experimentation, and the probabilistic attention mechanism of (zhou2023informer, ). This section is organised as follows. In Section 2.1 we present the background of time series forecasting using neural networks. In Section 2.2 we briefly recapitulate the canonical attention mechanism introduced in (vaswani2017attention, ). In Section 2.3 we summarise the concept of probabilistic attention and point out some of its weaknesses. Finally, in Section 2.4 we summarise the encoder-architecture of the vanilla transformer along with the positional encoding.
2.1 Time series forecasting using neural networks
Neural networks, particularly one-dimensional convolutional neural networks (1D CNNs) and Long Short-Term Memory (LSTM) networks, offer black-box solutions for time series forecasting (KhanZulfiqar, ; KimTaeYoung, ). 1D CNNs use convolutional layers to identify local patterns and hierarchies within time series data, but are unable to capture longer time dependencies (al2023lstm, ). Indeed, short-term prediction are greatly affected by locality, that is, the last few elements of the sequence greatly determine the trend of the model and, thus, the short-term prediction, unless seasonality and cyclic patterns play a role at that specific timestamp. LSTMs are able to take into account seasonality and cyclic patterns for medium-term dependencies thanks to the use memory cells, overcoming the vanishing gradient problem via gating mechanisms. However, the recursive nature of LSTMs and the memory cells limitations makes them less efficient and effective for long-term predictions (zhou2023informer, ).
Transformers have surpassed RNNs and LSTMs in several applications, such as NLP, thanks to their ability to model long-range dependencies, learn from large datasets, and parallelise computations efficiently. Motivated by these properties, transformers have been deemed as excellent candidates to the LSTF problem (li2019enhancing, ; zhou2023informer, ). Indeed, unlike traditional methods that rely on fixed time windows, lagged observations and/or recursive neural networks, transformers can process long sequences efficiently, making them suitable for both short-term and long-term forecasting tasks (lim2021temporal, ).
Transformers are complex encoder-decoder architectures that encode the sequence of samples and use the extracted information to predict the several future elements of the sequence via a decoder. The key element to understand these architectures is the attention mechanism, see (vaswani2017attention, ). In a nutshell, the main conclusion of (vaswani2017attention, ) is that performance in NLP applications can be significantly improved by eliminating the recurrence structure of deep learning models and instead focusing exclusively on exploiting the attention mechanism to extract relevant information from sequences of data. In this work we refer to this encoding-decoding model as the vanilla transformer.
When it comes to the application of transformers to time-series problems, progress has been significantly slower than in NLP. This is due to the fact that the attention mechanism presents two main limitations: loss of positional information, and computational and memory cost (li2019enhancing, ). The former problem is solved in the application of the vanilla transformer to NLP problems via a positional encoding of the input, and residual connections. However, the validity of this approach in the case of time series remains controversial, see, for instance (zeng2023transformers, ), suggesting that further research is needed. On the other hand, improving the computational and memory cost of the attention mechanism has actually been addressed in the time-series literature. The evaluation of the original attention mechanism has time and memory complexity $\Theta(n^{2})$, where $n$ is the width of this layer. As a consequence of the memory bottleneck, not every data scientist has access to the computing resources needed to train and evaluate a transformer model with the full attention mechanism on large sequences of data, preventing transformers from becoming a machine learning state of the art in time series forecasting in practice. The most remarkable solution to accelerate the attention mechanism for time series forecasting is that of (zhou2023informer, ), where the authors propose a probabilistic attention mechanism that runs in time and memory $\Theta(n\log n)$. Other proposed solutions to this problem involve applying other sorts of mechanisms to the sequence of data instead of the attention mechanism, see for instance (zhou2022fedformer, ).
@@ 2.2 Attention Mechanism
so the $i$-th output is a weighted average of the values. The scale factor $1/\sqrt{d_{q}}$ does not change the output of $\operatorname{softmax}$; its purpose is to help scaling the gradients of $\operatorname{softmax}$ during training in order to avoid gradients too close to 0, which may decrease the rate at which a transformer is able to learn from data.
The most common application of attention in transformers corresponds to the case $Q=K$. In this setting, the full attention mechanism resembles a smooth non-parametric regression (watson1964smooth, ; johnston1979smooth, ) of $V$ in terms of $Q=K$,  (nguyen2022fourierformer, ), effectively providing a smoothing of $V$. We refer to self-attention as the case when $Q=K=V$. Intuitively, self-attention weighs in the similarity of different elements in a sequence and smooths the input data. Empirically this process has been shown to be highly effective in capturing both local and global patterns in data (feng2024attention, ). However, this intuition has not been corroborated from a theoretical point of view, where mathematical results have only managed to unveil the Lipschitz constant of the attention mechanism (vuckovic2020mathematical, ).
@@
$\text{where }head_{i}$ $=\operatorname{Attn}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),$
@@
A critique to the full attention mechanism is the potential loss of positional information. First of all, note that the output of $\operatorname{FullAttn(Q,K,V)}$ is a matrix whose $i$-th row is $\sum_{j=1}a_{i,j}V_{j}$. This average $\sum_{j=1}a_{i,j}V_{j}$ results in a partial loss of position as it may be the case that $a_{i,j}$ is large for a $V_{j}$ that is far away from $V_{i}$ in the input sequence. That is, we lose the sense of temporal locality of the values as we average values with entries far away in the input sequence. We formalise this loss of local information is in Proposition 1, whose proof can be found in Appendix A. First, let us introduce some notation. For a permutation $\pi\colon\{0,1,\ldots,n-1\}\to\{0,1,\ldots,n-1\}$, the permutation matrix $P_{\pi}$ is the matrix with dimensions $n\times n$ defined as $P_{\pi}[i,j]=1$ for $j=\pi(i)$ and $P_{\pi}[i,j]=0$ otherwise.
@@
As we will see in Section 2.4, usually only the last layer of the vanilla transformer is a fully-linear layer; the rest of layers are attention layers with projections acting on the dimension of the characteristics. Therefore, positional information is only exploited in these last fully-linear layers. However, during the repeated application of attention layers positional information is reduced due to the averages carried out in the full attention mechanism. This issue is addressed by the implementation of the positional encoding, which aims to minimise this reduction of positional information when applying attention layers. However, it seems that, in the case of time series, positional encoding is not enough. Indeed, in experiments, on average a vanilla transformer trained for a particular time series produces similar predictions on an original input and a permuted version of this input  (zeng2023transformers, ).
@@ 2.3 Probabilistic Attention
Among state-of-the-art models, Informer is distinguished as one of the most frequently cited and top-performing (zhou2023informer, ). The authors propose a probabilistic attention mechanism that runs in time and memory $\Theta(n\log n)$. The key idea here is selecting $u$ queries, where $u=5\lceil\log n\rceil$, and setting the rest of the queries to zeroes. To select these queries, the authors introduce a sparsity measurement for the queries, which approximates the Kullback-Leibler divergence of the attention scores of a query and the uniform distribution. Finding the queries whose attention scores have largest sparsity measurement would require to compute the whole attention matrix to begin with. Instead, the authors propose a heuristic method: for each query $Q_{i}$, they sample $\Theta(\log n)$ scalar products of the form $\langle Q_{i},K_{j}\rangle$ and approximate the sparsity measurement of the query $Q_{i}$ using these samples. Let us denote by $\operatorname{ProbSelect}(Q,K)$ the method that selects the top $u$ queries of $Q$ with this sampling heuristic and masks the rest of the queries to zeroes. Then, the probabilistic attention mechanism is
@@
Even though the ProbAttn method does significantly reduce the time and memory complexity of the attention mechanism and seems to be able to capture the most significant attention scores in practice according to the experiments carried out in (zhou2023informer, ), it does lead to a significant loss of information. Indeed, the output of $\operatorname{ProbAttn(Q,K,V)}$ has $n-u$ entries that are equal to the average of the values. We believe this to be a major drawback of this attention mechanism, and this is corroborated in our experimentation. The authors of (zhou2023informer, ) try to fix this issue by applying residual connections in their model, as well as a complex architecture that exploits convolutions in conjunction with ProbAttn. However, our experiments show that this is not enough, as the vanilla transformer is able to significantly outperform Informer.
2.4 The Encoder-Decoder architecture and Positional Encoding
The encoder-decoder architecture in transformers forms the backbone of many advanced models in natural language processing and beyond. In this architecture, the encoder processes the input sequence, transforming it into a rich, contextual representation. Here we briefly describe the architecture proposed in (vaswani2017attention, ), which we use in our model in conjunction with LAM. Each encoder layer consists of a multi-head self-attention mechanism, which allows the model to weigh the importance of different parts of the input relative to each other, and two linear projections acting on the dimension of the variables (i.e. the last dimension of the input) with activation functions. Formally, these linear projections act as the matrix $W^{O}$ in the definition of $\operatorname{MultiHead}$, they are represented as $d_{model}\times d_{model}$ matrices, but they have LeakyReLu activation functions on the outputs. The purpose of projection layers is to learn a domain representation that boosts the performance of the attention mechanism.